# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v2

@@ Abstract
Large Language Models, though successful in code generation, struggle with code quality analysis because they are limited by static training data and can’t easily adapt to evolving best practices. We introduce MetaLint, an instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications. Unlike conventional approaches that train models on static code quality conventions, MetaLint employs instruction tuning on synthetic linter-generated data with dynamic conventions to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining. To evaluate this, we construct a benchmark of challenging idioms inspired by real-world coding standards such as Python Enhancement Proposals (PEPs) and assess whether MetaLint-trained models reason adaptively or simply memorize. Our results show that MetaLint training improves generalization to unseen idioms. Qwen3-4B attains a 70.37% F-score on a manually curated and challenging PEP idiom detection benchmark, achieving the highest recall (70.43%) among all evaluated models. For localization, it reaches 26.73%, which is a strong outcome for its 4B parameter size and comparable to larger state-of-the-art models such as o3-mini, highlighting its potential for future-proof code quality analysis. Furthermore, MetaLint training enables generalization in idiom detection across model families, model scales, synthetic data from diverse linters, and Java idioms, demonstrating the general applicability of our approach. We plan to release our code and data to enable reproducibility and further work.
@@ 1 Introduction
With the rise of Large Language Models (LLM) of code, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent Singhal et al. (2024); Zheng et al. (2024). Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Singhal et al., 2024; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023). Several evaluation studies indicate that LLMs struggle with this task Singhal et al. (2024); Zheng et al. (2024), while attempts to improve them through prompting or training are limited by task-specific, static datasets often grounded in narrow or outdated coding practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024; Zhang et al., 2024b). As a result, these systems often perform poorly when detecting rare issue types or when applied to code distributions that differ from their training data (Holden & Kahani, 2024). They may also over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024). Ideally, we would develop LLM systems that can identify code quality issues without explicit supervision for target idioms—especially hard or rare patterns—and adapt to evolving best practices over time.
We approach this problem by training the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms. For example, a commonly used idiom for generating secrets or passwords in Python is to use the random.choice standard library function. However, as noted in PEP 506 (D'Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against using this module for security reasons, which is often missed by developers, as highlighted by accepted answers on forums like StackOverflow. PEP 506 also introduces a more secure semantic block or idiom in the form of the secrets module and the secrets.choice function, which acts as a safer alternative to the random.choice idiom. As illustrated by this example, detecting and locating idioms associated with bad practices can be leveraged for identifying code quality issues like code smells (Wikipedia contributors, 2024) or Common Weakness Enumerations (CWE) (MITRE Corporation, 2024). Additionally, these issues can be addressed by replacing instances of ``bad'' idioms with corresponding ``good'' idioms that align with best practices. Moreover, for this example and similar abstract idioms, constructing a precise rule-based approach is difficult. Simply flagging any use of random.choice, even in non-security-critical scenarios (e.g., randomization in a game engine), could result in a poor user experience. Vijayvergiya et al. (2024) show that LLMs can capture abstract notions of code quality, such as code idioms where building a linter or rule-based approach is challenging, by incorporating semantic reasoning about code and developer intent.
In this work, we train LLMs to recognize code idioms through a higher-level instruction-following task dubbed ``meta-linting'': given a specification of a best-practice code idiom $I$, the model learns to identify and localize non-idiomatic code fragments. Our pipeline is designed to support easy-to-hard generalization (Sun et al., 2024b). The easy cases involve simple idioms that can already be captured by existing linters, while the hard cases correspond to nuanced patterns such as PEP 506, where constructing precise rule-based checks is infeasible. To enable this, we generate synthetic training data for easy idioms using available linters and leverage it to improve performance on harder cases where linter support is lacking. While prior work such as Zhang et al. (2024c; b) has explored automated refactoring of non-idiomatic Python code, including the use of LLMs with prompting, our focus differs in three ways. First, we target challenging idioms beyond the reach of current linters. Second, we train on easy idioms with the goal of transferring detection ability to harder cases. Finally, we emphasize adaptability, aiming for LLMs that can accommodate evolving best practices provided in-context as instructions and examples, rather than memorizing a static rule sets.
@@
To tackle meta-linting, we introduce MetaLint, a training framework motivated by prior work showing that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022). Since meta-linting treats each idiom as a distinct task or code quality judgment, instruction fine-tuning (IFT) and preference optimization (PO) naturally extend detection ability to novel idioms. Existing linters (e.g., Ruff (ruf, ) for Python and PMD (pmd, ) for Java) provide large-scale synthetic data by enforcing simple idioms, which we use both for supervised IFT and as verifiers during PO to improve performance on harder idioms. To systematically study this generalization, we construct a benchmark of challenging idioms derived from popular PEPs introducing high-level constructs. We evaluate state-of-the-art reasoning and code models on this benchmark and compare them with MetaLint trained models, examining whether they can move beyond memorizing easy idioms.
@@
    We introduce MetaLint, a training framework that leverages instruction following and synthetic data to enable easy-to-hard generalization while remaining adaptable to evolving best practices.
@@
    We construct a benchmark of challenging, broadly relevant code-quality idioms inspired by PEPs to evaluate the extent of easy-to-hard generalization achieved by MetaLint.
@@
    We benchmark state-of-the-art code and reasoning models on our PEP hard-idiom benchmark and compare them against MetaLint-trained models. Our method achieves the highest detection recall and competitive localization scores, even with smaller 4B models and without test-time compute.
@@
    We show that MetaLint generalizes across programming languages (Python, Java), model families (Qwen, Llama), linters (Ruff, PMD, Tree-Sitter), test-time reasoning settings (with and without CoT), and model scales (3B–8B).
@@ 2 Related Work
Code Quality Analysis with Large Language Models. A large body of prior work has explored the use of LLMs for code quality analysis through code review and static analysis. Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as rule-guided linters via prompting or fine-tuning. While Blyth et al. (2025) proposes a static analysis-driven prompting framework to improve LLM-generated code, Du et al. (2025) conversely uses LLMs to enhance static analysis tools by reducing false-positives. LintLLM (Fang et al., 2025) and (Shin et al., 2025) leverages LLMs for linting of Verilog and Quantum computing code. Khare et al. (2023) show LLMs outperform traditional static analysis tools for security-related CWEs with step-by-step reasoning. Vijayvergiya et al. (2024) train LLMs for best practice violation detection and localization, while Rasheed et al. (2024) design a multi-agent review pipeline for maintainability, efficiency, and bugs. Other works (Jiang et al., 2025b; Yao et al., 2025) use prefix-tuning and reinforcement learning with static analysis–based rewards for higher-quality, functionally correct code generation. Naik et al. (2024) and Jaoua et al. (2025) integrate LLMs with linters to produce more informative code reviews. RIdiom (Zhang et al., 2024c) introduces a rule-based way to identify and refactor non-idiomatic Python code with AST rewrite rules, while Zhang et al. (2024b) combines LLMs and rule-based detectors but doesn't explore nuanced idioms like PEP 506 or training LLMs to keep up with evolving best practices. Finally, CoUpJava (Jiang et al., 2025a) presents Java version upgrade benchmarks, conceptually similar to our hard PEP idiom benchmark for Python. Although prior work demonstrates the potential of LLMs for code quality tasks, it focuses on fixed rule sets or best practices that require retraining as they evolve. In contrast, we train models to interpret high-level specifications and perform static analysis, enabling broader generalization.
Instruction Following for Generalization. Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning fixed tasks. Prior work shows diverse task instructions allow models to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022). Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020). Instructions serve as high-density task representations, substituting supervision (Puri et al., 2022) and enabling generalization even with minimal labeled data or pseudo-labeled examples (Gu et al., 2022). Studies also show that instruction diversity drives generalization, with varied instructions outperforming repeated exposure to identical formats (Charton et al., 2024). This phenomenon holds across domains, including program synthesis where task-level prompting facilitates generalization in code generation models (Niu et al., 2023). SELF-GUIDE (Zhao et al., 2024) performs task-specific instruction following using synthetic data, demonstrating effectiveness, but relying entirely on LLM-generated data without verifiers. These results suggest instruction tuning acts as task-level meta-learning, enabling models to adapt to new tasks through natural language. Building on this we model specific code quality idioms as individual tasks and generate large-scale synthetic data for each meta-task to support cross-idiom generalization. This allows the trained model to keep pace with new idioms and evolving best practices. We also discuss additional related work on easy-to-hard generalization in Appendix B.
@@ 3 Method
We design the MetaLint framework to teach an LLM to operationalize idiom descriptions provided in context, rather than memorizing specific idioms, thereby enabling adaptation to novel idioms at test time. We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}=\{D_{I},E_{I}\}$. The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$. This setup discourages rote memorization and encourages adaptive reasoning over the prompt’s specification, since flagging violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$. By framing best practices as meta-tasks, this approach enables the LLM to remain flexible and better aligned with evolving best practices. We describe the components of our training framework in Figure 1 and Figure 2 below.
@@ 3.1 Synthetic Data Generation
One of the main goals of our meta-task formulation is enabling easy-to-hard generalization. We train LLMs on a set of ``easy'' idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by a linter). Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand and detect code idioms from in-context descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to the untrained model. Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision. For Python, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, etc., while for Java, we use the PMD static analyzer, which covers 269 idioms as well as some manually written tree-sitter [^1] https://tree-sitter.github.io/tree-sitter/ queries inspired by 8 Java Enhancement Protocols (JEPs) (Table 8). We run Ruff, PMD, and the JEP tree-sitter queries on Python and Java source code files $f\in\mathcal{F}$ from the STACK (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories. This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$. Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data. Additionally, to automatically build the meta-task instruction prompts $M_{I_{\mathcal{L}}}$ for each idiom, we scrape rule-specific documentation from the Ruff and PMD websites, including descriptions and examples. An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix C.1. For the JEP tree-sitter queries, since they are few in number, we manually write the meta-task prompts.

### v3

@@ Abstract
Large Language Models excel at code generation but struggle with code quality analysis, where best practices evolve and cannot be fully captured by static training data. We introduce MetaLint, a training framework that treats code quality analysis as detecting best practice violations from high-level specifications over semantic code fragments (code idioms). Instead of training on a fixed set of rules, MetaLint reorganizes supervision around dynamically specified best practices using synthetic linter-derived labels, integrated with instruction-following and preference optimization. This encourages extrapolation to more complex, unseen best practices at test time, consistent with easy-to-hard generalization without retraining. To evaluate MetaLint, we create a new benchmark of hard-to-detect best practices inspired by Python Enhancement Proposals. Across this benchmark, MetaLint improves generalization to unseen best practices. Qwen3-4B achieves a 2.7× detection F-score gain (25.9% → 70.4%), the highest recall, and a 26.7% localization F-score, matching larger models such as o3-mini. These gains generalize across programming languages, model families, scales, reasoning settings, and linter sources.
Easy to Hard Generalization, Transfer Learning
@@ 1 Introduction
Code quality analysis offers a unique opportunity to study generalization in Large Language Models (LLMs) because best practices are contextual, evolving, and poorly captured by static training data. With the rise of LLMs for code, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent (Singhal et al., 2024; Zheng et al., 2024). Researchers have explored LLMs’ potential to improve code quality through benchmarks, code review agents, and static analysis (Singhal et al., 2024; Vijayvergiya et al., 2024; Fang et al., 2025). However, studies show that LLMs often struggle with these tasks, and improvements via prompting or training are limited by static datasets reflecting narrow or outdated best practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024; Zhang et al., 2024b). As a result, these systems often perform poorly on rare best practices, different code distributions (Holden & Kahani, 2024) or over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024). This difficulty stems not only from limited or outdated training data, but also from supervision: models are trained on a fixed set of best-practice labels, which encourages memorization of individual rules rather than learning reusable abstractions. More broadly, generalization to harder or novel violations depends less on model scale or optimization and more on how the learning problem is structured (Chang & Bisk, 2025).
Motivated by this view, we frame code quality analysis as meta-learning (or ``meta-linting'') over code idioms, recurring patterns that capture semantic intent beyond surface syntax. In this formulation, best practices are semantic patterns over idioms, and violations occur when problematic idioms appear in contexts where safer or more appropriate alternatives exist (see Appendix C). We operationalize this by using linters to generate synthetic supervision for idioms, then train LLMs to detect and locate violations from high-level natural language specifications and illustrative examples (see Figure 1). Code idioms provide a natural intermediate abstraction, bridging simple syntactic cues and full program semantics. This allows models to leverage simpler idioms to support detection of more complex, context-dependent best practices. For example, in Python, generating passwords or tokens using random.choice is common but insecure in security-sensitive contexts. PEP 506 (D'Aprano, 2017) recommends using secrets.choice instead. Detecting such violations requires conditional reasoning: not every random.choice is problematic, but its use in security-critical code constitutes a violation. Our approach trains models to recognize both the idiom (random.choice) and the relevant context, allowing them to flag violations that go beyond simple syntactic patterns.
This highlights a broader learning challenge. Some best practices are relatively easy to detect because they rely on surface-level cues. For example, existing linters such as Ruff include rules (e.g., S104–S108) that flag likely hard-coded secrets using heuristic signals like credential-shaped strings or suspicious variable names. While limited and incomplete, these rules expose recurring cues associated with security-sensitive code. Meanwhile, hard best practices, such as those captured by PEP 506, build on these same cues but require additional semantic and contextual reasoning. Reliably detecting insecure randomness involves recognizing when such signals co-occur with developer intent, such as credential generation or authentication logic, going beyond what fixed heuristics can capture.
This defines a difficulty progression in code quality analysis driven by the increasing contextual and semantic reasoning required to recognize best practice violations. As observed by Chang & Bisk (2025), failures to generalize along such progressions stem not from data scarcity, model scale, or inference limits, but from learning paradigms that fix the hypothesis class and inhibit abstraction. In code quality analysis, flat supervision maps programs directly to violation labels, forcing models to memorize isolated best practices rather than extract reusable structure. Although easy violations expose low-level heuristics, current training setups do not allow these signals to be elevated into higher-order representations, preventing inductive transfer to harder or novel cases. In contrast, MetaLint restructures supervision around code idioms, using signals from easy violations to learn semantic abstractions that become the basis for detecting harder best practice violations from high-level specifications.
@@
MetaLint integrates naturally with existing instruction fine-tuning (IFT) and preference-optimization (PO) methods, using them for exploiting task structure to achieve cross-task generalization (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022). Existing linters (e.g., Ruff (ruf, ) for Python and PMD (pmd, ) for Java) provide large-scale synthetic data by enforcing easy best practices, which we use both for supervised IFT and as verifiers during PO to improve performance on hard best practices. To systematically study this generalization, we manually construct a benchmark of hard best practices derived from popular PEPs requiring contextual reasoning. We evaluate state-of-the-art reasoning and code models on this benchmark and compare them with MetaLint trained models, examining whether they can move beyond memorizing easy heuristics.
@@
    We introduce MetaLint, a training framework that reorganizes supervision to bootstrap learning signals for easy best practices from linters to support inductive transfer to hard and novel best practices.
@@
    We construct a benchmark of hard-to-detect, broadly relevant best practices inspired by PEPs to probe whether MetaLint models exhibit easy-to-hard generalization.
@@
    We benchmark state-of-the-art code and reasoning models on our PEP hard best practice benchmark and compare them against MetaLint-trained models. With our training Qwen3-4B outperforms larger models like o3-mini achieving the highest detection recall and comparable localization scores without test-time compute.
@@
    We show that MetaLint generalizes across programming languages (Python, Java), model families (Qwen, Llama), linters (Ruff, PMD, Tree-Sitter), reasoning settings (with and without CoT), and model scales (3B–8B).
Overall, MetaLint is not a replacement for linters, but an empirical demonstration that organizing supervision around semantic units enables LLMs to transfer learned code quality reasoning to harder or novel best practices. Finally, code quality analysis offers a realistic testbed for studying generalization: while correctness is largely objective, the difficulty progression spans from cases verifiable by automated linters to harder, context-dependent cases requiring human judgment (PEP benchmark). This work takes a first step toward such settings, where supervision becomes increasingly incomplete along the difficulty progression.
@@ 2 Related Work
Code Quality Analysis with Large Language Models. A large body of prior work has explored using LLMs for code quality analysis by interfacing them with static analysis tools or fixed rule sets. Travis Fischer (2024); lin (2023) treat LLMs as rule-guided linters via prompting or fine-tuning. While Blyth et al. (2025) proposes a static analysis-driven prompting framework to improve LLM-generated code, Du et al. (2025) conversely uses LLMs to enhance static analysis tools by reducing false-positives. Fang et al. (2025); Shin et al. (2025); Khare et al. (2023) leverage LLMs for linting and show they can outperform traditional static analysis tools. Vijayvergiya et al. (2024) train LLMs for best practice violation detection and localization, while Rasheed et al. (2024) design a multi-agent review pipeline for code review. Other works (Jiang et al., 2025b; Yao et al., 2025) use prefix-tuning and reinforcement learning with static analysis–based rewards for higher-quality, functionally correct code generation. Naik et al. (2024); Kapadnis et al. (2025); Jaoua et al. (2025) provide linter results to LLMs for more informative code reviews. Zhang et al. (2024c, b) explore AST rewrite rules and hybrid approaches combining LLMs and rules for code quality refactoring. Finally, CoUpJava (Jiang et al., 2025a) propose a Java version upgrade benchmark, mined using AST refactoring rules. Across these approaches, static analysis tools and handcrafted rules typically act as either the final decision mechanism or the primary training signal, anchoring model behavior to a fixed set of best practices. In contrast, we use static analysis as supervision for the easy end of a difficulty progression, and train models to detect best practice violations from high-level descriptions, enabling transfer to harder and previously unseen cases that extend beyond linter coverage.
Easy-to-Hard Generalization. Recent work suggests that easy-to-hard generalization arises not merely from curriculum ordering, but from how supervision is structured to promote abstraction across related tasks (Chang & Bisk, 2025). In mathematical reasoning, models trained on easier problems (e.g., levels 1–3) consistently generalize better to harder benchmarks (e.g., levels 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025), with several studies emphasizing the role of high-quality supervision, particularly for harder instances (He et al., 2024). Sun et al. (2024) show that reward models trained on simple code and math problems improve performance on more complex tasks. More broadly, (Hu et al., 2025; Gaunt et al., 2016) demonstrate that training on structurally simpler instances can yield robustness to longer or more complex reasoning problems, including code. Similarly, Zhang et al. (2024a) evaluate reward models on algorithmic tasks such as string manipulation and observe transfer from simpler to more complex formats. Most closely, Chang & Bisk (2025) argue that easy-to-hard generalization is primarily driven by the learning paradigm and finding the right hypothesis class, rather than data scarcity, model scale or inference limits. While prior work studies this phenomenon in domains such as counting, mathematics, or contest coding, code quality analysis offers a complementary setting where correctness is largely objective but supervision becomes increasingly incomplete along the difficulty progression. Additional related work on leveraging instruction following for generalization is discussed in Appendix D.
@@ 3 Method
We design the MetaLint framework to reorganize supervision for best practice violation detection, shifting the hypothesis class from memorizing fixed rules to identifying violations from high-level natural language descriptions of problematic code idioms, in line with recent work showing that such reorganization is critical for easy-to-hard generalization (Chang & Bisk, 2025). We formulate best practice violation detection as an instruction-following meta-task $M_{I}$, where for a given problematic idiom $I$, the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}=\{D_{I},E_{I}\}$. The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$. This setup discourages rote memorization by construction, as correctness is defined relative to the provided specification, and encourages adaptive reasoning over the prompt’s description. For example flagging best practice violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$. By framing best practices as meta-tasks, this approach supports the LLM in remaining flexible and better aligned with evolving best practices. We describe the components of our training framework in Figure 1 and Figure 2 below.
@@ 3.1 Synthetic Data Generation
One of the main goals of our meta-task formulation is to change the learning task to detecting best practice violatons from high level specifications to encourage transfer from easy to hard cases. We train LLMs on a set of ``easy'' best practice idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by a linter). Our hypothesis is that training on $I_{\mathcal{L}}$ exposes the model to a structured family of best practice idiom specifications, which supports inductive transfer to hard to detect best practice idioms in $I_{\mathcal{L^{\prime}}}$. Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision. For Python, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, etc., while for Java, we use the PMD static analyzer, which covers 269 idioms as well as some manually written tree-sitter [^1] https://tree-sitter.github.io/tree-sitter/ queries inspired by 8 Java Enhancement Protocols (JEPs) (Table 8). We run Ruff, PMD, and the JEP tree-sitter queries on Python and Java source code files $f\in\mathcal{F}$ from the STACK (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories. This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$. Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data. Additionally, to automatically build the meta-task instruction prompts $M_{I_{\mathcal{L}}}$ for each idiom, we scrape rule-specific documentation from the Ruff and PMD websites, including descriptions and examples. An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix E.1. For the JEP tree-sitter queries, since they are few in number, we manually write the meta-task prompts.